How much energy is needed to change a 2 kg block of -8.2 Celsius ice to 10.6 Celsius water
Answer #1

S_water = 4186 J/kg oC

specific heat of ice = S_ice = 2100 J/kg oC


Lice = latent heat of melting of ice = 3.34*10^5 J/kg

ice at -8.2 is converted to ice at 0 celsius


ice at 0oC is converted to water at 0 celsius


water at 0oC is converted to water at 10.6 celsius


energyneeded = m*sice*(0 - Tice) + m*Lice + m*Swater(Tf - 0)


energy neede = 2*2100*(0 - (-8.2) ) + 2*3.34*10^5 + 2*4186*(10.6-0) = 791183.2 J